A 95 kg patient swallows a 34 ?Ci beta emitter whose half-life is 5.0 days and whose RBE is 1.6. The beta particles are emitted with an average energy of 0.35 MeV, 90% of which is absorbed by the body.

You are a health care worker needing to find the patient's dose equivalent after one week. These series of steps will help you find that dose equivalent. In all questions, assume the radioactive nuclei are distributed throughout the patient's body and are not being excreted

Part A: How many radioactive nuclei did the patient swallow?  

N0 = 7.84x10^11

Part A: How many radioactive nuclei did the patient swallow?  

N = 2.97x10^11

Part C

How many radioactive decays occured in the patient during the first week?  
Answer #1

Az 34 uci = 34610-6Ci =(34X10-9 (3.7x100) =1258x1dfeq;ty2 = Sder ca Rot 1/2 (1* 258X106) (5 X 24X60X60) ena 7. 848 10 No= Rb) N = No (0.5t/t1/2 > 7.84XLO (0.5) 715 N = 2.97% 10HRadioactive nuclei occurred in week first NO -N =No (1-(1) 75) = 7.84X10(1-10)53) ND = 4.878 holl